A product is designed to have length of 20.cm+/ 0.1 cm. The output of the manufacturing process is identified to be centered at 20.0155 cm and the standard deviation is estimated at .069 cm. Determine the capability of the system Cp

A product has a target length of 20.0 cm, and the process has a standard value for the standard deviation of 0.05 cm. Calculate the standard deviation control chart upper limit when the sample size is 5"

A product is designed to have length of 20.cm+/ 0.1 cm. The output of the manufacturing process is identified to be centered at 20.0155 cm and the standard deviation is estimated at .069 cm. Determine proportion of noncomforing product produced by this process
Answer #1

USL ao LSL = 19.9 bo009 USL LSL 0 2 - 0.489 6x0.069 99 Mt36 UCL= ao+ 3x0. o5 =20 + 0,06子 = 2D-06子チ し0. - P/D19.1- 20.0155 X-2